OpenStudy (anonymous):

A town has a population of 1000 in 1990. It has been decreasing by 3% ever since. Make an exponential function model with x the years since 1990 and y the population

OpenStudy (anonymous):

y=ab^x y=1000(1-0.03)^x y=1000(0.97)^x